# FSGe: A fast and strongly-coupled 3D fluid-solid-growth interaction method

### v1

@@ Abstract
Equilibrated fluid-solid-growth (FSGe) is a fast, open source, three-dimensional (3D) computational platform for simulating interactions between instantaneous hemodynamics and long-term vessel wall adaptation through growth and remodeling (G&R). Such models are crucial for capturing adaptations in health and disease and following clinical interventions. In traditional G&R models, this feedback is modeled through highly simplified fluid models, neglecting local variations in blood pressure and wall shear stress (WSS). FSGe overcomes these inherent limitations by strongly coupling the 3D Navier-Stokes equations for blood flow with a 3D equilibrated constrained mixture model (CMMe) for vascular tissue G&R. CMMe allows one to predict long-term evolved mechanobiological equilibria from an original homeostatic state at a computational cost equivalent to that of a standard hyperelastic material model. In illustrative computational examples, we focus on the development of a stable aortic aneurysm in a mouse model to highlight key differences in growth patterns and fluid-solid feedback between FSGe and solid-only G&R models. We show that FSGe is especially important in blood vessels with asymmetric stimuli. Simulation results reveal greater local variation in fluid-derived WSS than in intramural stress (IMS). Thus, differences between FSGe and G&R models became more pronounced with the growing influence of WSS relative to pressure. Future applications in highly localized disease processes, such as for lesion formation in atherosclerosis, can now include spatial and temporal variations of WSS.
@@ 1 Introduction
The intricate interaction between hemodynamics and growth (change in mass) and remodeling (change in microstructure) of a blood vessel plays a vital role in development, homeostasis, and disease progression [1]. The constrained mixture model (CMM) simulates the continuous deposition and degradation of tissue constituents [2, 3], often based on the concept of mechanical homeostasis, which is visualized in Figure 1. In this model of mechanobiological homeostasis, changes in vascular composition and properties stem from deviations in regulated variables from defined set points. These regulated variables primarily include intramural stress (IMS), $\sigma_{I}$, and wall shear stress (WSS), $\tau_{w}$, which have consistently been found to be maintained at homeostatic values under physiologic conditions in vivo [4].
Hemodynamics $\mathcal{F}$ are a crucial determinant of the biomechanical state of a blood vessel, characterized in part by its regulated variables. Solid mechanics $\mathcal{S}$ within the vessel wall define blood vessel geometry and microstructure. Changes in the solid mechanics thus trigger changes in hemodynamics, $\mathcal{F}$, rendering the feedback loop a fluid-solid-growth interaction (FSG) problem. Most prior work considered either constant or reduced-order hemodynamics. The novelty of our work is a strongly coupled feedback between hemodynamics $\mathcal{F}$ and solid mechanics $\mathcal{S}$ in a fully three-dimensional (3D) framework based on the equilibrated constrained mixture model (CMMe) [5]. In the following, we review previous approaches in more detail and highlight the contributions of this work.
@@
Hemodynamics in vivo typically include complex flow phenomena such as flow separation and recirculation induced by changes in cross-sectional area or bifurcations [6]. Consequently, pressure and WSS typically show a high spatial and temporal variation. Local hemodynamics can be quantified with computational fluid dynamics (CFD) or fluid-structure interaction (FSI) solutions [7]. However, the CMM commonly includes simplified assumptions for axial pressure gradient $\Delta p$ and WSS $\tau_{w}$ based on Hagen–Poiseuille law [5]
$\Delta p=\frac{8\mu lQ}{\pi r^{4}},\quad\tau_{w}=\frac{4\mu Q}{\pi r^{3}},$ (1)
@@
with dynamic blood viscosity $\mu$, vessel length $l$ and radius $r$, and bulk blood flow rate $Q$. These equations assume fully developed steady-state flow through a straight cylindrical tube with axisymmetric flow and zero radial and circumferential components, which is generally not upheld in in vivo settings.
Reference Fluid Solid Stimuli Coupling Geometry
@@
Figueroa et al. [8] 3D Navier-Stokes Membrane CMM pulsatile mean weak idealized
@@
Sheidaei et al. [9] 3D Navier-Stokes Membrane CMM pulsatile mean weak in vivo
@@
Watton et al. [10] 3D Navier-Stokes 3D hybrid pulsatile mean weak in vivo
@@
Grytsan et al. [11] 3D Navier-Stokes 3D hybrid steady-state weak in vivo
@@
Latorre et al. [12] Control-volume 3D CMMe steady-state strong idealized
Schwarz et al. [13] 3D Navier-Stokes 3D CMM steady-state strong in vivo
@@
The local interaction between hemodynamics and vessel growth and remodeling (G&R) mechanisms is crucial in numerous diseases and their treatments, as, for example, in tissue-engineered vascular grafts (TEVGs) [14, 15], vein graft adaptation and failure following coronary artery bypass graft surgery [16, 17], and in congenital heart patients with pulmonary disease [18, 19]. For an overview of G&R models, i.e., kinematic, constrained mixture model (CMM), and hybrid, see [20]. A non-comprehensive overview of similar fluid-solid-growth interaction models is given in Table 1. Figueroa et al. [8] developed a method that couples three-dimensional (3D) hemodynamics to a membrane model of the vessel wall. This theory is based on the theory of small on large [21], superimposing "small" deformations during the cardiac cycle upon "large" deformations due to G&R. The membrane wall formulation is related to the coupled momentum method for fluid-structure interaction (FSI) in blood vessels [22]. The long-term G&R deformation is coupled to pulsatile hemodynamics, illustrated in an example of aneurysmal development. This approach was used in a patient-specific aneurysm geometry in Sheidaei et al. [9]. The work of Latorre et al. [12] couples 3D G&R with a simplified control-volume formulation of fluid dynamics. Pressure and wall shear stress (WSS) are derived from an axial node-based control-volume approach based on the Hagen-Poiseuille law (1) with additional pressure loss terms. For a more detailed discussion of reduced-order fluid dynamics, see Pfaller et al. [23]. Most recently, Schwarz et al. [13] coupled 3D CMM and 3D hemodynamics. While this constitutes the most detailed approach so far, it comes at high computational cost due to the need to integrate heredity integrals point-wise in full CMM.
Most FSG approaches reviewed above, except for Latorre et al. [12] and Schwarz et al. [13], employ weak coupling between fluid and solid G&R, also known as loose or weak coupling. Here, the fluid and solid domains are solved independently, with their solutions being exchanged only once in each load step. This approach assumes that the interaction between the fluid and solid is relatively weak, allowing for some degree of decoupling between the domains. Loose coupling is typically used for problems where the fluid and solid have a minor influence on each other or when computational efficiency is prioritized over solution accuracy. On the other hand, strong coupling schemes rely on an iterative procedure that ensures the convergence of fluid and solid solutions.
This work introduces an equilibrated fluid-solid-growth (FSGe) method to strongly couple 3D G&R based on the equilibrated constrained mixture model (CMMe) with 3D hemodynamics. Using CMMe greatly reduces the computational cost of evaluating the solid model and generally arrives at the desired model prediction in much fewer steps. In the following, we introduce the governing differential equations of the fluid and solid domains (Section 2). We then review computational techniques employed to solve the coupled fluid-solid problem (Section 3). In illustrative computational examples of an asymmetric aneurysm, we compare key differences between a solid-only G&R model and our FSGe formulation (Section 4). We close with a discussion of the results, limitations, and future perspectives (Section 5).
@@ 2 Governing equations of fluid and solid domains
As outlined in Figure 1, the FSGe model relies on the solution of fluid (2.1) and solid (2.2) mechanics, whose governing equations we briefly review in this section. Additionally, for completeness we introduce both the full CMM (2.2.1) and CMMe (2.2.2) G&R models. Both 3D fluid and solid were solved numerically in our open-source multi-physics solver svFSIplus [24], the C++ version of svFSI [25]; both solvers are released with the SimVascular project [26].
@@ 2.2 Solid growth and remodeling
We follow the classic approach of finite strain theory to model the vessel wall displacement, $\bm{d}=\bm{x}-\bm{X}$, with reference position $\bm{X}$ and current position $\bm{x}$. We calculate the deformation gradient $\bm{F}$, the Jacobian $J$, and the right Cauchy–Green tensor $\bm{C}$ as
@@
The balance of linear momentum in the case of negligible inertial loads [29] in the solid domain $\Omega^{\text{s}}$ yields the quasi-static problem
@@
with solid stress tensor $\bm{\sigma}^{\text{s}}$ provided constitutively by the G&R material and its solid time scale $T^{\text{s}}$. We prescribe boundary conditions
@@
We use a G&R constitutive law to relate deformation $\bm{F}$ to stresses $\bm{\sigma}^{\text{f}}$. In the following, we briefly review the full CMM and the CMMe. The CMM has been diversely applied, e.g., in altered flow and pressure conditions [31], abdominal [32] and thoracic [33] aortic aneurysms, vein grafts [34], or tissue-engineered vascular grafts [35]. However, due to its computational complexity, there are only limited implementations in three dimensions [36, 37]. In this work, we use the fast, rate-independent CMMe method that directly predicts the long-term evolved solution that is in mechanobiological equilibrium [38]. This material model had been previously implemented in the open-source finite element solver FEBio [39], with computational costs comparable to hyperelastic materials [5]. Prior work relied on constant-pressure models [5, 33, 40] and control-volume hemodynamics [12]. We briefly introduce the CMM in Section 2.2.1 and outline the CMMe in Section 2.2.2 but refer interested readers to the original publications.
Figure 2: Blood vessel with length $l$, radius $r$, and thickness $h$. Its constituents are elastin $e$ (isotropic), smooth muscle $m$ (red, circumferential), and four collagen fiber families $c_{i}$ (circumferential, axial, diagonal angle $\alpha_{0}$). Image created by Sebastian L. Fuchs and licensed under the Creative Commons Attribution 4.0 International License.
@@ 2.2.1 Constrained Mixture Model
The CMM assumes that structurally significant constituents $\alpha$ with individual reference configurations, material properties, mass balance relations, and turnover rates are constrained to move together. The mass production per unit volume at G&R time $\tau\in[0,s]$, where $s$ is the current G&R time and $\tau$ is an intermediate time at which constituents are deposited, is expressed as
@@
$\Upsilon^{\alpha}(\tau)=1+K^{\alpha}_{\sigma}\,\Delta\sigma_{I}(\tau)-K^{\alpha}_{\tau}\,\Delta\tau_{w}(\tau_{w}),$ (5)
@@
$\Delta\sigma_{I}(\tau)=\frac{\sigma_{I}(\tau)}{\sigma_{Io}}-1,\quad\Delta\tau_{w}(\tau)=\frac{\tau_{w}(\tau)}{\tau_{wo}}-1.$ (6)
governed by changes in intramural stress (IMS) $\sigma_{I}$ (induced by pressure and axial force) and WSS $\tau_{W}$ (induced by blood flow) from homeostatic set points $\sigma_{Io}$ and $\tau_{wo}$, respectively. The constituent removal between deposition times $\tau$ and current time $s$ is modeled by an exponential decay
$q^{\alpha}(s,\tau)=\exp\left(-\int_{\tau}^{s}k^{\alpha}(t)\,\mathrm{d}t\right),\quad k^{\alpha}(t)=k^{\alpha}_{o}\,\left[1+K^{\alpha}\,\Delta\sigma^{2}(t)\right],$
where $k^{\alpha}$ is a rate parameter with homeostatic value $k^{\alpha}_{o}$, and $K^{\alpha}>0$ is a gain parameter for IMS deviation. We can express the mass density per current volume of each constituent $\rho^{\alpha}_{R}$ by
$\rho^{\alpha}(s)=\int_{-\infty}^{s}m^{\alpha}(\tau)\,q^{\alpha}(s,\tau)\,\mathrm{d}\tau,$ (7)
where $m^{\alpha}$ is the mass density production rate. The constituent-specific elastic strain energy density functions $\hat{W}^{\alpha}$ for elastin $e$, collagen $c$, and smooth muscle $m$ are characterized by the neo-Hookean and Fung-type exponential relations as
@@
$W(s)=\sum_{\alpha}W^{\alpha}(s),\quad W^{\alpha}(s)=\int_{-\infty}^{s}\frac{m^{\alpha}(\tau)\,q^{\alpha}(s,\tau)}{\rho(s)}\,\hat{W}^{\alpha}_{n(\tau)}(s)\,\mathrm{d}\tau.$

### v2

@@ Abstract
Equilibrated fluid-solid-growth (FSGe) is a fast, open source, three-dimensional (3D) computational platform for simulating interactions between instantaneous hemodynamics and long-term vessel wall adaptation through mechanobiologically equilibrated growth and remodeling (G&R). Such models can capture evolving geometry, composition, and material properties in health and disease and following clinical interventions. In traditional G&R models, this feedback is modeled through highly simplified fluid solutions, neglecting local variations in blood pressure and wall shear stress (WSS). FSGe overcomes these inherent limitations by strongly coupling the 3D Navier-Stokes equations for blood flow with a 3D equilibrated constrained mixture model (CMMe) for vascular tissue G&R. CMMe allows one to predict long-term evolved mechanobiological equilibria from an original homeostatic state at a computational cost equivalent to that of a standard hyperelastic material model. In illustrative computational examples, we focus on the development of a stable aortic aneurysm in a mouse model to highlight key differences in growth patterns between FSGe and solid-only G&R models. We show that FSGe is especially important in blood vessels with asymmetric stimuli. Simulation results reveal greater local variation in fluid-derived WSS than in intramural stress (IMS). Thus, differences between FSGe and G&R models became more pronounced with the growing influence of WSS relative to pressure. Future applications in highly localized disease processes, such as for lesion formation in atherosclerosis, can now include spatial and temporal variations of WSS.
@@ 1 Introduction
The intricate interaction between hemodynamics and growth (change in mass) and remodeling (change in microstructure) of a blood vessel plays a vital role in development, homeostasis, and disease progression [1]. The constrained mixture model (CMM) simulates the continuous deposition and degradation of tissue constituents [2, 3], often based on the concept of mechanical homeostasis, which is visualized in Figure 1. In this model of mechanobiological homeostasis, changes in vascular composition and properties stem from deviations in regulated variables from defined set points. These regulated variables primarily include scalar metric measures of IMS, $\sigma_{I}$, and WSS, $\tau_{w}$, which have consistently been found to be maintained near homeostatic values under physiologic conditions in vivo [4].
Hemodynamics are a crucial determinant of the biomechanical state of a blood vessel, characterized in part by its regulated variables. Solid mechanics within the vessel wall define blood vessel geometry and microstructure. Changes in the solid mechanics thus trigger changes in hemodynamics rendering this feedback a fluid-solid-growth interaction (FSG) problem. Most prior work considered either constant or reduced-order hemodynamics. The novelty of our work is a strongly coupled feedback between hemodynamics and solid mechanics in a fully 3D framework based on the mechanobiologically CMMe [5]. In the following, we review previous approaches in more detail and highlight the contributions of this work.
@@
Hemodynamics in vivo typically include complex flow phenomena such as separation and recirculation induced by changes in cross-sectional area or bifurcations [6]. Consequently, pressure and WSS typically show a high spatial and temporal variation. Local hemodynamics can be quantified with computational fluid dynamics (CFD) or fluid-structure interaction (FSI) solutions [7]. However, the CMM commonly includes simplified assumptions for axial pressure gradient $\Delta p$ and WSS $\tau_{w}$ based on Hagen-Poiseuille law [5]
$\Delta p=\frac{8\mu lQ}{\pi{\color[rgb]{0,0,0}a}^{4}},\quad\tau_{w}=\frac{4\mu Q}{\pi{\color[rgb]{0,0,0}a}^{3}},$ (1)
@@
with dynamic blood viscosity $\mu$, vessel length $l$ and inner radius $a$, and bulk blood flow rate $Q$. These equations assume fully developed steady-state flow through a straight cylindrical tube with axisymmetric flow and zero radial and circumferential components, which is generally not upheld in in vivo settings.
The local interaction between hemodynamics and vessel G&R mechanisms is crucial in numerous diseases and their treatments, as, for example, in aneurysms [8, 9, 10, 11], tissue-engineered vascular grafts [12, 13], vein graft adaptation and failure following coronary artery bypass graft surgery [14, 15], and in congenital heart patients with pulmonary disease [16, 17]. For an overview of G&R models, i.e., kinematic [18], CMM [2], and rate-based constrained mixture model (rb-CMM) [19], see [20]. A non-comprehensive overview of similar fluid-solid-growth interaction models is given in Table 1.
Baek et al. [21, 22] devised a simple control-volume network model for the flow that was coupled to a membrane CMM. Figueroa et al. [23] developed a method that couples 3D hemodynamics to a membrane model of the vessel wall. This theory is based on the theory of small on large [24], superimposing "small" deformations during the cardiac cycle upon "large" deformations due to G&R. The membrane wall formulation is related to the coupled momentum method for FSI in blood vessels [25]. The long-term G&R deformation is coupled to pulsatile hemodynamics, illustrated in an example of aneurysmal development. This approach was used in a patient-specific aneurysm geometry in Sheidaei et al. [26].
Watton et al. (2009) [8] integrated the membrane model from Watton et al. (2004) [19] into an FSG framework, simulating saccular aneurysms on idealized 3D arterial geometries. Aparicio et al. [9] considered cyclic deformations as a G&R stimulus for collagen with a spatially and temporally heterogeneous endothelium. Selimovic et al. [10] showed that spatially heterogeneous WSS can account for asymmetry in intracranial aneurysms. Teixeria et al. [11] considered the wall shear stress aspect ratio [27] as a G&R stimulus, a metric for oscillatory flow.
The work of Latorre et al. [28] couples 3D G&R with a simplified control-volume formulation of fluid dynamics. Pressure and WSS are derived from an axial node-based control-volume approach based on the Hagen-Poiseuille law (1) with additional pressure loss terms. For a more detailed discussion of reduced-order fluid dynamics, see Pfaller et al. [29]. Most recently, Schwarz et al. [30] coupled 3D CMM and 3D hemodynamics. While this constitutes the most detailed approach so far, it comes at high computational cost due to the need to integrate heredity integrals point-wise in full CMM.
Most FSG approaches reviewed above, except for Baek et al. [22], Latorre et al. [28], and Schwarz et al. [30], employ weak coupling between fluid and solid G&R. That is, the fluid and solid domains are solved independently, with their solutions being exchanged only once in each load step. This approach assumes that the interaction between the fluid and solid is relatively weak, allowing for some degree of decoupling between the domains. Loose coupling is typically used for problems where the fluid and solid have a minor influence on each other or when computational efficiency is prioritized over solution accuracy. On the other hand, strong coupling schemes rely on an iterative procedure that ensures the convergence of fluid and solid solutions.
Reference Fluid Solid Stimuli FS-Coupling Geometry
Baek [21, 22] Control-volume membrane CMM pulsatile mean strong idealized
@@
Figueroa [23] 3D Navier-Stokes membrane CMM pulsatile mean weak idealized
@@
Sheidaei [26] 3D Navier-Stokes membrane CMM pulsatile mean weak in vivo
@@
Watton [8] 3D Navier-Stokes membrane rb-CMM pulsatile mean weak idealized
Aparicio [9] 3D Navier-Stokes membrane rb-CMM pulsatile mean, cyclic weak idealized
@@
Selimovic [10] 3D Navier-Stokes membrane rb-CMM steady-state weak in vivo
Grytsan [31] 3D Navier-Stokes 3D rb-CMM steady-state weak in vivo
Teixeria [11] 3D Navier-Stokes 3D rb-CMM pulsatile & oscillatory weak in vivo
@@
Latorre [28] Control-volume 3D CMMe steady-state strong idealized
Schwarz [30] 3D Navier-Stokes 3D CMM steady-state strong in vivo
@@
This work introduces an equilibrated fluid-solid-growth (FSGe) method to strongly couple 3D growth and remodeling (G&R) based on the equilibrated constrained mixture model (CMMe) with 3D hemodynamics. Using CMMe greatly reduces the computational cost of evaluating the solid model and generally arrives at the desired model prediction in much fewer steps. In the following, we introduce the governing differential equations of the fluid and solid domains (Section 2). We then review computational techniques employed to solve the coupled fluid-solid problem (Section 3). In illustrative computational examples of an asymmetric aneurysm, we compare key differences between a solid-only G&R model and our FSGe formulation (Section 4). We close with a discussion of the results, limitations, and future perspectives (Section 5).
@@ 2 Governing equations of fluid and solid domains
As outlined in Figure 1, the FSGe model relies on the solution of fluid (2.1) and solid (2.2) mechanics, whose governing equations we briefly review in this section. Additionally, for completeness we introduce both the full constrained mixture model (CMM) (2.2.1) and CMMe (2.2.2) G&R models. Both three-dimensional (3D) fluid and solid were solved numerically in our open-source multi-physics solver svFSIplus [32], the C++ version of svFSI [33]; both solvers are released with the SimVascular project [34].
@@ 2.2 Solid growth and remodeling
We follow the classic approach of finite strain theory to model the vessel wall displacement, $\bm{d}=\bm{x}-\bm{X}$, with reference position $\bm{X}$ and current position $\bm{x}$. We calculate the deformation gradient $\bm{F}$, the Jacobian $J$, and the right Cauchy-Green tensor $\bm{C}$ as
@@
The balance of linear momentum in the case of negligible body forces and inertial loads [37] in the solid domain $\Omega^{\text{s}}$ yields the quasi-static problem
@@
with solid stress tensor $\bm{\sigma}^{\text{s}}$ provided constitutively by the G&R constitutive relation and its solid time scale $T^{\text{s}}$. We prescribe boundary conditions
@@
We use a G&R constitutive law to relate deformation $\bm{F}$ to stresses $\bm{\sigma}^{\text{s}}$. In the following, we briefly review the full CMM and the CMMe. The CMM has been diversely applied, e.g., in altered flow and pressure conditions [39], abdominal [40] and thoracic [41] aortic aneurysms, vein grafts [42], or tissue-engineered vascular grafts [43]. However, due to its computational complexity, there are only limited implementations in three dimensions [44, 45]. In this work, we use the fast, rate-independent CMMe method that directly predicts the long-term evolved solution that is in mechanobiological equilibrium [46]. This material model had been previously implemented in the open-source finite element solver FEBio [47], with computational costs comparable to hyperelastic materials [5]. Prior work relied on constant-pressure models [5, 41, 48] and control-volume hemodynamics [28]. We briefly introduce the CMM in Section 2.2.1 and outline the CMMe in Section 2.2.2 but refer interested readers to the original publications.
Figure 2: Blood vessel with length $l$, inner radius $a$, and thickness $h$. Its constituents are elastin $e$ (isotropic), smooth muscle $m$ (red, circumferential), and four collagen fiber families $c_{i}$ (circumferential, axial, diagonal angle $\alpha_{0}$). Image created by Sebastian L. Fuchs and licensed under the Creative Commons Attribution 4.0 International License.
@@ 2.2.1 Constrained Mixture Model
The CMM assumes that structurally significant constituents $\alpha$ with individual reference configurations, material properties, mass balance relations, and turnover rates are locally constrained to move together. The mass production per unit volume at G&R time $\tau\in[0,s]$, where $s$ is the current G&R time and $\tau$ is an intermediate time at which constituents are deposited, is expressed as
@@
$\Upsilon^{\alpha}(\tau)=1+K^{\alpha}_{\sigma}\,\Delta\sigma_{I}(\tau)-K^{\alpha}_{\tau}\,\Delta\tau_{w}(\tau),$ (5)
@@
$\Delta\sigma_{I}(\tau)=\frac{\sigma_{I}(\tau)}{\sigma_{Io}}-1,\quad\Delta\tau_{w}(\tau)=\frac{\tau_{w}(\tau)}{\tau_{wo}}-1,$ (6)
governed by changes in the scalar metrics intramural stress (IMS) $\sigma_{I}$ (induced by pressure and axial force) and wall shear stress (WSS) $\tau_{W}$ (induced by blood flow) from homeostatic set points $\sigma_{Io}$ and $\tau_{wo}$, respectively. The constituent removal between deposition times $\tau$ and current time $s$ is modeled by an exponential decay
$q^{\alpha}(s,\tau)=\exp\left(-\int_{\tau}^{s}k^{\alpha}(t)\,\mathrm{d}t\right),\quad k^{\alpha}(t)=k^{\alpha}_{o}\,\left[1+K^{\alpha}\,\Delta\sigma_{I}^{2}(t)\right],$
where $k^{\alpha}$ is a rate parameter with homeostatic value $k^{\alpha}_{o}$, and $K^{\alpha}>0$ is a gain parameter for IMS deviation. We can express the mass density per current volume of each constituent $\rho^{\alpha}$ by
$\rho^{\alpha}(s)=\int_{-\infty}^{s}m^{\alpha}(\tau)\,q^{\alpha}(s,\tau)\,\mathrm{d}\tau.$ (7)
The constituent-specific elastic strain energy density functions $\hat{W}^{\alpha}$ for elastin $e$, collagen $c$, and smooth muscle $m$ are characterized by the neo-Hookean and Fung-type exponential relations as
@@
$W(s)=\sum_{\alpha}W^{\alpha}(s),\quad W^{\alpha}(s)=\int_{-\infty}^{s}\frac{m^{\alpha}(\tau)\,q^{\alpha}(s,\tau)}{\rho(s)}\,\hat{W}^{\alpha}_{n(\tau)}\left(\bm{F}^{\alpha}_{n(\tau)}(s)\right)\,\mathrm{d}\tau.$